A stove has a surface temperature of 500 K. The room it is in has a temperature of 310 K. The emissivity of the stove is 0.9, and its surface area is 4.5 m2. What is the net power radiated into the room by the stove?
1
Expert's answer
2020-04-28T09:47:21-0400
"P=e\sigma A(T_2^4-T_1^4)"

"P=(0.9)( 5.67 \cdot 10^{−8} )(4.5)(500^4-310^4)\\=12231\ W"